OpenStudy (anonymous):

x=3y-4 a) write the equation in slope-inter form b)give the slop and y-intercept of the line

OpenStudy (agreene):

So, slope-intercept form is y=mx+b where m is your slope and b is your y-intercept That means that first you have to solve for y. x=3y-4 add 4 to each side x+4 = 3y divide each side by 3 (x+4)/3 = y you can distribute the division to: x/3 + 4/3 = y and if you wish, convert x/3 to 1/3 x so you see the slope easier y= 1/3x + 4/3 That makes 1/3 =m = slope and 4/3 = b = y-intercept :)